OpenStudy (anonymous):

Help please? If f is the antiderivative of (x^2/(1+x^5)) such that f(1) = 0 then f(4) =?

OpenStudy (anonymous):

I understand this is integration but i do not know how to do it sadly:(

OpenStudy (anonymous):

\[\int\limits_{1}^{4} \frac{ x^2 }{ 1+x^5 } dx\] I guess this is what you have to integrate

OpenStudy (anonymous):

ooh so the 1 and 4 are the starting and stopping values... but how do you integrate it?

OpenStudy (anonymous):

Haha, well it's a bit complicated

OpenStudy (anonymous):

Is there a general formula i can follow?

OpenStudy (anonymous):

You can try checking to see if \(x^5+1\) has any real roots, and maybe get some partial fractions going...

OpenStudy (anonymous):

Siths is much better at this stuff than me haha, I'll let him take over :P

OpenStudy (anonymous):

Ok! Sith would the real roots be like x = -1 and x = -1^2/5?

OpenStudy (anonymous):

\[x^5+1=(x+1)(x^4-x^3+x^2-x+1)\] so then \[\begin{align*}\frac{x^2}{x^5+1}&=\frac{A}{x+1}+\frac{Bx^3+Cx^2+Dx+E}{x^4-x^3+x^2-x+1}\\\\ x^2&=(A+B)x^4+(-AB+C)x^3+(A+C+D)x^2\\ &\quad+(-A+D+E)x+(A+E) \end{align*}\]

OpenStudy (anonymous):

There's only one real root to consider here, the other four are complex.

OpenStudy (anonymous):

I think we can approximate it with simpsons rule as well?

OpenStudy (anonymous):

@iambatman we're not computing a definite integral here, but rather solving a differential equation. The given info, that \(f(1)=0\), is an initial condition. Once we compute the antiderivative, we have an unknown constant, i.e. when we get a solution like \[f(x)=(\text{some expression in terms of }x)+C\] we plug in the initial condition and get \[0=(\text{expression with }x\text{ replaced with }1\text{s})+C\] and we can solve for \(C\).

OpenStudy (anonymous):

Ah, gotcha :)

OpenStudy (anonymous):

ok so i understand the whole general expression + c but i do not know why you did the previous breakdown of the equation with the ABCs

OpenStudy (anonymous):

It's the partial fraction decomposition. We have some rational function of the form \(\dfrac{P(x)}{Q(x)}\), where \(P\) and \(Q\) are polynomials. We break up \(Q(x)\) into irreducible factors \(S_1(x),\ldots,S_n(x)\) (also polynomials) such that \[\frac{P(x)}{Q(x)}=\sum_{k=1}^n\frac{R_k(x)}{S_k(x)}\] where \(R_k(x)\) is one degree less than each corresponding \(S_k(x)\). Basically what this means is that we can split up a rational function into more convenient rational functions. It's the reverse of the process shown here, for instance: \[\frac{x}{x^2+1}+\frac{1}{x-1}=\frac{x(x-1)+(x^2+1)}{(x^2+1)(x-1)}=\frac{2x^2-x+1}{x^3-x^2+x-1}\] The \(A,B,C\ldots\) are unknown coefficients that we want to find so that \[\text{complicated rational function}=\text{sum of less complicated rational functions}\]

OpenStudy (anonymous):

oooh I see, so now that the equation has been broken up, how would one find each coefficient since it equals to x^2? I vaguely understand the reverse process but im kind of lost on when it comes to finding each letter

OpenStudy (anonymous):

on the equation p(x)/q(x) does that work to find every integral?

OpenStudy (anonymous):

Here's a nice resource for review: http://tutorial.math.lamar.edu/Classes/CalcII/PartialFractions.aspx For this particular decomposition, it's a matter of solving a system of 5 equations with 5 unknowns, namely the one you get from matching up the coefficients of corresponding powers of \(x\). For instance, if I had \[x=(A+B)x^2+(B-A)x\] then I get the system \[\begin{cases}A+B=0\\B-A=1\end{cases}\] For this function, you have \[0x^4+0x^3+x^2+0x+0=(A+B)x^4+(-A+B+C)x^3+(A+C+D)x^2\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+(-A+D+E)x+(A+E)\] which gives the system, \[\begin{cases}\begin{align*}A+B&=0\\-A+B+C&=0\\A+C+D&=1\\-A+D+E&=0\\A+E&=0\end{align*}\end{cases}\]

OpenStudy (anonymous):

oooh, i find the link VERY helpful, do you mind if i go and try this on my own? Your explanation is great as well btw!

OpenStudy (anonymous):

By all means. And you're welcome!